# Explaining Tournament Solutions with Minimal Supports

### v2

@@ Appendix A Proofs for Section 4
As this proof follows the same idea used in the proof of Proposition 1, we only provide a sketch of the proof. Let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be a MS for $w\in\operatorname{wUC}(G)$. If exists $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)<\lceil(n+1)/2\rceil$ or for all $c^{\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}}(w,c^{\prime})+\mu_{\mathcal{X}}(c^{\prime},c)<n+1$, we prove there exists a completion $\mathcal{X}^{\prime}$ of $\mathcal{X}$ such that none of the conditions of Lemma 7 is satisfied for $c$. Hence, $w\not\in\operatorname{MM}(\mathcal{X}^{\prime})$, $w\not\in\text{NW}_{\operatorname{MM}}(\mathcal{X})$ and $\mathcal{X}$ is not a MS for $w\in\operatorname{wUC}(G)$. Contradiction. ∎
@@
The inclusion in NP is clear. To prove NP-hardness, we proceed with a reduction from the NP-complete set cover problem (Karp 1972). Given a set of $p$ element $U=\{e_{1},e_{2},\dots,e_{n}\}$ and a set $S$ of $q$ subsets of $U$, $S=\{s_{1},s_{2},\dots,s_{m}\}$ such that $\bigcup_{s\in S}s=U$, and an integer $k$, the set cover problem looks at the existence of a subset $D\subseteq S$ of size at most $k$ such that $\bigcup_{s\in D}s=U$. With this notation, we now prove that there exists a set cover of size less than $k\in\mathbb{N}$ if and only if there exists a MS of size less than $p+q+k$ for a specific tournament.
Let $U$, $S$, $k$ be an instance of set cover. First, we can safely assume that there exists no $s\in S$ such that $S=U$. We define a complete 2-weighted tournament $G(U,S)$ ($G$ for short) with the following $p+q+1$ candidates: a candidate $w$ that is the winner for whom we compute a SMS, $p$ candidates $c_{e_{j}}$, one for each elements of $U$, and $q$ candidates $c_{s_{j}}$ for the elements of $S$. We define the weight function $\mu$ according to the adjacency matrix described in Table 2.
@@
$c_{e_{1}}$
$c_{e_{n}}$
2 1
if $e_{i}\in s_{j}$
then 1
else 2
$c_{s_{1}}$
$c_{s_{m}}$
0
if $e_{j}\in s_{i}$
then 1
else 0
1
@@
We now show that if there exists a set cover of $(U,S)$ of size less than $k$ then there exists a MS of $G(U,S)$ of size less than $p+q+k$ Suppose there exists a set cover $D$ of $(U,S)$ of size less than $k$. Let $\mathcal{X}\subseteq G$ such that: $\forall s\in D$, $\mu_{\mathcal{X}}(w,s)=2$, $\forall s\in S\setminus{D}$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},s)=1$ and $\forall e\in U$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},e)=1$. $|\mathcal{X}|=2|D|+1(p+q-|D|)=p+q+k$. $\forall s\in D$, $\mu_{\mathcal{X}}(w,c_{s})=2\geq\left\lceil(2+1)/2\right\rceil$. $\forall s\in S\setminus{D}$, $\exists s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(c_{s^{\prime}},c_{s})+\mu_{\mathcal{X}}(w,c_{s^{\prime}})=1+2\geq 2+1$. $\forall e\in U$, since $D$ is a cover set of (U,S), $\exists s\in D$ such that $\mu_{\mathcal{X}}(c_{s},c_{e})+\mu_{\mathcal{X}}(w,c_{s})=1+2\geq 2+1$. Hence, $w\in\text{NW}_{\operatorname{wUC}}(\mathcal{X})$ and there exists a MS contained in $\mathcal{X}$, i.e., of size at most $p+q+k$.
We now show the converse, if there exists a MS of $G(U,S)$ of size less than $p+q+k$ then there exists a set cover of $(U,S)$ of size less than $k$. Suppose there exists a MS $\mathcal{X}$ for $w\in\operatorname{wUC}(G)$ such that $|\mathcal{X}|=p+q+k$. Let $D=\{c:c\in\mathcal{C}\setminus{\{w\}},\mu(w,c)\geq\left\lceil(2+1)/2\right\rceil\}$ be the set of dominion of $w$. By minimality of a MS, $\forall c\in D$, $\forall c^{\prime}\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(c^{\prime},c)=0$ and $\forall c\in\mathcal{C}\setminus{D}$, $\exists!c^{\prime}\in D$, $\mu_{\mathcal{X}}(c^{\prime},c)=1$ and $\forall c^{\prime\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}}(c^{\prime\prime},c)=0$. Hence, p+q+k=2|D|+1(p+q|D|)=p+q+|D|p+q+k=2|D|+1(p+q-|D|)=p+q+|D| and $|D|=k$. However, since $\mu_{X}\leq\mu$ and $\forall e\in U$, $\mu(w,c_{e})=0$, $\forall e\in U$, $\exists c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu_{X}(c,c_{e})=1$ and $\mu_{X}(w,c)=2$. Additionally, $\forall e\in U$, $\forall c\in\mathcal{C}$, $\mu(c,c_{e})\geq 1$ and $\mu(w,c)\geq 2$ if and only if $\exists s\in S$ such that $c=c_{s}$ and $e\in s$. Thus, $\forall e\in U$, $\exists s\in S$ such that $e\in s$ and $s\in D$, i.e., $D$ is a set cover of $(U,S)$ of size $k$. ∎
@@
We start with the lower bound. Let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be a MS for $w\in\operatorname{wUC}(G)$, Thanks to Proposition 8, we know that for all losing candidates $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)\geq\lceil(n+1)/2\rceil$ or exists a unique $c^{\prime}$ such that $\mu_{\mathcal{X}}(w,c^{\prime})+\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1$. In particular, we have $\max_{c^{\prime\prime}\neq w}\mu_{\mathcal{X}}(w,c^{\prime\prime})+\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1$ which implies $\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1-\max_{c^{\prime\prime}\neq w}\mu_{\mathcal{X}}(w,c^{\prime\prime})$. Let $D=\{c:c\in\mathcal{C},\mu_{\mathcal{X}}(w,c)\geq\lceil(n+1)/2\rceil\}$ and $k=|D|$. If $k=0$,
@@
Note that $n+1-\max_{c\neq w}\mu_{\mathcal{X}}(w,c)\geq 1$ and since, $\max_{c\neq w}\mu_{\mathcal{X}}(w,c)\geq\left\lceil(n+1)/2\right\rceil$, $\max_{c\neq w}\mu_{\mathcal{X}}(w,c)+\lceil(n+1)/2\rceil-n-1\geq n+1$. Hence, $|\mathcal{X}|\geq n+m-2$.
We prove that the bound is tight. Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a candidate $w\in\mathcal{C}$ such that there exist a candidate $c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu(w,c)=n$ and for all candidates $c^{\prime}\in\mathcal{C}\setminus{\{w,c\}}$, $\mu(c,c^{\prime})\geq 1$. For all $c^{\prime}\in\mathcal{C}\setminus{\{w\}}$, exists $c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu(w,c)+\mu(c,c^{\prime})=n+1$. Hence, $w\in\operatorname{wUC}(G)$. Then let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ such that $\mu(w,c)=n$ and for all candidates $c^{\prime}\in\mathcal{C}\setminus{\{w,c\}}$, $\mu(c,c^{\prime})=1$ and $\mu_{X}=0$ everywhere else. $|\mathcal{X}|=n+m-2$ and for all completions $\mathcal{X}^{\prime}$ of $\mathcal{X}$, $\mu_{\mathcal{X}^{\prime}}(w,c)=n\geq\lceil(n+1)/2\rceil$ and for all candidates $c^{\prime}\in\mathcal{C}\setminus{\{w,c\}}$, $\mu_{\mathcal{X}^{\prime}}(w,c^{\prime})+\mu_{\mathcal{X}^{\prime}}(c,c^{\prime\prime})=n+1$.
For the upper bound, thanks to Proposition 8, we know that $n+1$ pairwise comparisons are enough to prove that $w$ cover one losing candidate. This gives us a global upper bound of (m1)(n+1)(m-1)(n+1) which is tight. Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a candidate $w\in\mathcal{C}$ such that for all candidate $c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu(w,c)=1$ and exists a unique candidate $d_{c}$ such that $\mu(d_{c},c)=0$. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, for all $c^{\prime}\in\mathcal{C}\setminus{\{w,d_{c}\}}$, $\mu(w,c)<\lceil(n+1)/2\rceil$, $\mu(w,c)+\mu(c,c^{\prime})=\mu(w,c)+n-\mu(c^{\prime},c)\leq 1+n-1<n+1$ and $\mu(w,c)+\mu(c,d_{c})=\mu(w,c)+n-\mu(d_{c},c)=n+1$. Hence, $w\in\operatorname{wUC}(G)$ and the only possible MS $\mathcal{X}=(\mathcal{C},\mu_{X})$ is for all candidate $c\in\mathcal{C}\setminus{\{w\}}$, $\mu(w,c)=1$, $\mu(d_{c},c)=n$ and $\mu_{X}=0$ everywhere else. $|\mathcal{X}|=(m-1)(1+n)$ and since it is the only MS, it is also a SMS. ∎
A.2 Proofs for Section 4.3
@@
Building a MS for $w\in\operatorname{MM}(G)$ is rather straightforward. We simply need to ensure that $w$ wins a minimum number $t$ of pairwise comparisons in each head-to-head with another candidate and ensure that each losing candidate loses at least ntn-t pairwise comparisons against another candidate. Suppose we do not do that and let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be a MS for $w\in\operatorname{MM}(G)$. Suppose for all $t\leq\left\lceil n/2\right\rceil$ exists $c_{0}\in\mathcal{C}\setminus{\{w\}}$ such that $\mu_{\mathcal{X}}(w,c_{0})<t$ and for all $c^{\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}}(c^{\prime},c_{0})<n-t$. Then let $\mathcal{X}^{\prime}$ be a completion of $\mathcal{X}$ such that $\mu_{\mathcal{X}^{\prime}}(w,c_{0})=\mu_{\mathcal{X}}(w,c_{0})$ and for all $c^{\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}^{\prime}}(c^{\prime},c_{0})=\mu_{\mathcal{X}}(c^{\prime},c_{0})$. Then the maximin score of $w$ is $\sigma_{w}\operatorname{MM}(w,\mathcal{X}^{\prime})\geq\mu_{\mathcal{X}^{\prime}}(w,c_{0})<t$ while the maximin score of $c_{0}$ is $\sigma_{c_{0}}\operatorname{MM}(c_{0},\mathcal{X}^{\prime})\geq n-\max_{c^{\prime}\in\mathcal{C}}\mu_{\mathcal{X}}(c^{\prime},c_{0})>n-(n-t)>t$. Since $\sigma_{\operatorname{MM}}(c_{0},\mathcal{X}^{\prime})>\sigma_{\operatorname{MM}}(w,\mathcal{X}^{\prime})$, $w\not\in\text{NW}_{\operatorname{MM}}(\mathcal{X})$. Contradiction. ∎
@@
The lower bound is achieved by maximizing the values of $t$ and $k=|\{c:c\in\mathcal{C},\,\mu(w,c)\geq n-t\}|$. Typically, if for all candidate $c\in\mathcal{C}\setminus{\{w\}}$, $\mu(w,c)=n$, we have $t=\left\lfloor n/2\right\rfloor$ and k=m1k=m-1 and $|\mathcal{X}|=(n-\left\lfloor n/2\right\rfloor)(m-1)+n(m-1-(m-1)))=\left\lceil n/2\right\rceil(m-1)$. Conversely, if $k=0$, $|\mathcal{X}|=(n-t)0+n(m-1-0))=n(m-1)$. ∎
@@
Following the ideas developed in Proposition 12, we introduce Algorithm 1. The SMS returned by the algorithm maximizes the maximin score of $w$, unless its score is larger than $\left\lfloor n/2\right\rfloor$, because ensuring a larger score would break the subset minimality of the SMSs. Hence, the maxwin property is satified. ∎

### v3

@@ Appendix A Proofs for Section 4
See 2
Proof.
Let $\mathcal{X}=(C,E_{\mathcal{X}})$ be an MS for $w\in\operatorname{TC}(G)$. Let us now show that every candidate different from $w$ has an in-degree of 1 in $\mathcal{X}$ and that $w$ has an in-degree of 0. According to Proposition 1, for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a directed path $p_{c}$ from $w$ to $c$ in $\mathcal{X}$. Suppose that there exists $c^{*}\in\mathcal{C}\setminus{\{w\}}$ such that there exist distinct $c^{\prime},c^{\prime\prime}$ such that $\{(c^{\prime},c^{*}),(c^{\prime\prime},c^{*})\}\subseteq E_{\mathcal{X}}$. Then $p_{c^{\prime}}\circ(c^{\prime},c^{*})$ and $p_{c^{\prime\prime}}\circ(c^{\prime\prime},c^{*})$ are two distinct directed paths from $w$ to $c$ in $\mathcal{X}$. Contradiction. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, every candidate different from $w$ has an in-degree of 1 in $\mathcal{X}$. Suppose that there exists $c^{*}\in\mathcal{C}\setminus{\{w\}}$ such that $(c^{*},w)\in E_{\mathcal{X}}$. No $p_{c}$ contains $(c^{*},w)$ otherwise as they start at $w$, they would pass by $w$ twice. Let $\mathcal{Y}=(\mathcal{C},E_{\mathcal{Y}})$ be such that $E_{\mathcal{Y}}=E_{\mathcal{X}}\setminus{\{(c^{*},w)\}}$. For all $c\in\mathcal{C}\setminus{\{w\}}$, $p_{c}$ is a directed path from $w$ to $c$ in $\mathcal{Y}$. Thus $w\in\text{NW}_{\operatorname{TC}}(\mathcal{Y})$ but $\mathcal{Y}\subsetneq\mathcal{X}$ so $\mathcal{X}$ is not an MS for $w\in\operatorname{TC}(G)$. Contradiction. Hence, $w$ has an in-degree of 0 in $\mathcal{X}$. Finally, since the sum of the in-degree of all nodes is the number of edges in a graph, $|\mathcal{X}|=m-1$. ∎
See 3
Proof.
With Dijkstra’s algorithm (dljkstra1959note), it is possible to build directed-rooted tree in $G$ with $w$ in $\mathcal{O}(|E|+|C|)$, i.e., $\mathcal{O}(m^{2})$. The shortest-path property is ensured by the fact that Dijkstra’s algorithm minimizes the cost of the directed path to every candidate. (Since we are working with unweighted graphs, simply exploring the graph in breadth first search will work). Hence, for every candidate $c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu(w,c)=1$, $c$ will be a child of $w$ in the obtained directed-rooted tree. ∎
A.2 Proofs for Section 4.2
@@
As this proof follows the same idea used in the proof of Proposition 1, we only provide a sketch of the proof. Let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be an MS for $w\in\operatorname{wUC}(G)$. If exists $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)<\lceil(n+1)/2\rceil$ or for all $c^{\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}}(w,c^{\prime})+\mu_{\mathcal{X}}(c^{\prime},c)<n+1$, we prove there exists a completion $\mathcal{X}^{\prime}$ of $\mathcal{X}$ such that none of the conditions of Lemma 7 is satisfied for $c$. Hence, $w\not\in\operatorname{MM}(\mathcal{X}^{\prime})$, $w\not\in\text{NW}_{\operatorname{MM}}(\mathcal{X})$ and $\mathcal{X}$ is not an MS for $w\in\operatorname{wUC}(G)$. Contradiction. ∎
@@
The inclusion in NP is clear. To prove NP-hardness, we proceed with a reduction from the NP-complete Set Cover problem (karp1972reducibility). Given a set of $p$ element $U=\{e_{1},e_{2},\dots,e_{n}\}$ and a set $S$ of $q$ subsets of $U$, $S=\{s_{1},s_{2},\dots,s_{m}\}$ such that $\bigcup_{s\in S}s=U$, and an integer $k$, the set cover problem looks at the existence of a subset $D\subseteq S$ of size at most $k$ such that $\bigcup_{s\in D}s=U$. With this notation, we now prove that there exists a set cover of size less than $k\in\mathbb{N}$ if and only if there exists an MS of size less than $p+q+k$ for a specific tournament.
Let $U$, $S$, $k$ be an instance of set cover. First, we can safely assume that there exists no $s\in S$ such that $S=U$. We define a complete 2-weighted tournament $G(U,S)$ ($G$ for short) with the following $p+q+1$ candidates: a candidate $w$ that is the winner for whom we compute an SMS, $p$ candidates $c_{e_{j}}$, one for each element of $U$, and $q$ candidates $c_{s_{j}}$ for the elements of $S$. We define the weight function $\mu$ according to the adjacency matrix described in Table 2.
@@
$c_{e_{1}}$ $c_{e_{n}}$ 2 1 if $e_{i}\in s_{j}$ then 1 else 2
$c_{s_{1}}$ $c_{s_{m}}$ 0 if $e_{j}\in s_{i}$ then 1 else 0 1
@@
We now show that if there exists a set cover of $(U,S)$ of size less than $k$ then there exists an MS of $G(U,S)$ of size less than $p+q+k$ Suppose there exists a set cover $D$ of $(U,S)$ of size less than $k$. Let $\mathcal{X}\subseteq G$ such that: $\forall s\in D$, $\mu_{\mathcal{X}}(w,s)=2$, $\forall s\in S\setminus{D}$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},s)=1$ and $\forall e\in U$, $\exists!s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(s^{\prime},e)=1$. $|\mathcal{X}|=2|D|+1(p+q-|D|)=p+q+k$. $\forall s\in D$, $\mu_{\mathcal{X}}(w,c_{s})=2\geq\left\lceil(2+1)/2\right\rceil$. $\forall s\in S\setminus{D}$, $\exists s^{\prime}\in D$ such that $\mu_{\mathcal{X}}(c_{s^{\prime}},c_{s})+\mu_{\mathcal{X}}(w,c_{s^{\prime}})=1+2\geq 2+1$. $\forall e\in U$, since $D$ is a cover set of (U,S), $\exists s\in D$ such that $\mu_{\mathcal{X}}(c_{s},c_{e})+\mu_{\mathcal{X}}(w,c_{s})=1+2\geq 2+1$. Hence, $w\in\text{NW}_{\operatorname{wUC}}(\mathcal{X})$ and there exists an MS contained in $\mathcal{X}$, i.e., of size at most $p+q+k$.
We now show the converse, if there exists an MS of $G(U,S)$ of size less than $p+q+k$ then there exists a set cover of $(U,S)$ of size less than $k$. Suppose there exists an MS $\mathcal{X}$ for $w\in\operatorname{wUC}(G)$ such that $|\mathcal{X}|=p+q+k$. Let $D=\{c:c\in\mathcal{C}\setminus{\{w\}},\mu(w,c)\geq\left\lceil(2+1)/2\right\rceil\}$ be the set of dominion of $w$. By minimality of an MS, $\forall c\in D$, $\forall c^{\prime}\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(c^{\prime},c)=0$ and $\forall c\in\mathcal{C}\setminus{D}$, $\exists!c^{\prime}\in D$, $\mu_{\mathcal{X}}(c^{\prime},c)=1$ and $\forall c^{\prime\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}}(c^{\prime\prime},c)=0$. Hence, p+q+k=2|D|+1(p+q|D|)=p+q+|D|p+q+k=2|D|+1(p+q-|D|)=p+q+|D| and $|D|=k$. However, since $\mu_{X}\leq\mu$ and $\forall e\in U$, $\mu(w,c_{e})=0$, $\forall e\in U$, $\exists c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu_{X}(c,c_{e})=1$ and $\mu_{X}(w,c)=2$. Additionally, $\forall e\in U$, $\forall c\in\mathcal{C}$, $\mu(c,c_{e})\geq 1$ and $\mu(w,c)\geq 2$ if and only if $\exists s\in S$ such that $c=c_{s}$ and $e\in s$. Thus, $\forall e\in U$, $\exists s\in S$ such that $e\in s$ and $s\in D$, i.e., $D$ is a set cover of $(U,S)$ of size $k$. ∎
@@
We start with the lower bound. Let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be an MS for $w\in\operatorname{wUC}(G)$, Thanks to Proposition 8, we know that for all losing candidates $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)\geq\lceil(n+1)/2\rceil$ or exists a unique $c^{\prime}$ such that $\mu_{\mathcal{X}}(w,c^{\prime})+\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1$. In particular, we have $\max_{c^{\prime\prime}\neq w}\mu_{\mathcal{X}}(w,c^{\prime\prime})+\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1$ which implies $\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1-\max_{c^{\prime\prime}\neq w}\mu_{\mathcal{X}}(w,c^{\prime\prime})$. Let $D=\{c:c\in\mathcal{C},\mu_{\mathcal{X}}(w,c)\geq\lceil(n+1)/2\rceil\}$ and $k=|D|$.
If $k=0$,
@@
Note that $n+1-\max_{c\neq w}\mu_{\mathcal{X}}(w,c)\geq 1$ and since, $\max_{c\neq w}\mu_{\mathcal{X}}(w,c)\geq\left\lceil(n+1)/2\right\rceil$, $\max_{c\neq w}\mu_{\mathcal{X}}(w,c)+\lceil(n+1)/2\rceil\geq n+1$. Hence, $|\mathcal{X}|\geq n+m-2$.
We prove that the bound is tight. Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a candidate $w\in\mathcal{C}$ such that there exist a candidate $c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu(w,c)=n$ and for every candidate $c^{\prime}\in\mathcal{C}\setminus{\{w,c\}}$, $\mu(c,c^{\prime})\geq 1$. For all $c^{\prime}\in\mathcal{C}\setminus{\{w\}}$, exists $c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu(w,c)+\mu(c,c^{\prime})=n+1$. Hence, $w\in\operatorname{wUC}(G)$. Then let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ such that $\mu(w,c)=n$ and for every candidate $c^{\prime}\in\mathcal{C}\setminus{\{w,c\}}$, $\mu(c,c^{\prime})=1$ and $\mu_{X}=0$ everywhere else. $|\mathcal{X}|=n+m-2$ and for all completions $\mathcal{X}^{\prime}$ of $\mathcal{X}$, $\mu_{\mathcal{X}^{\prime}}(w,c)=n\geq\lceil(n+1)/2\rceil$ and for every candidate $c^{\prime}\in\mathcal{C}\setminus{\{w,c\}}$, $\mu_{\mathcal{X}^{\prime}}(w,c^{\prime})+\mu_{\mathcal{X}^{\prime}}(c,c^{\prime\prime})=n+1$.
For the upper bound, thanks to Proposition 8, we know that $n+1$ pairwise comparisons are enough to prove that $w$ cover one losing candidate. This gives us a global upper bound of (m1)(n+1)(m-1)(n+1) which is tight. Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a candidate $w\in\mathcal{C}$ such that for every candidate $c\in\mathcal{C}\setminus{\{w\}}$ such that $\mu(w,c)=1$ and exists a unique candidate $d_{c}$ such that $\mu(d_{c},c)=0$. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, for all $c^{\prime}\in\mathcal{C}\setminus{\{w,d_{c}\}}$, $\mu(w,c)<\lceil(n+1)/2\rceil$, $\mu(w,c)+\mu(c,c^{\prime})=\mu(w,c)+n-\mu(c^{\prime},c)\leq 1+n-1<n+1$ and $\mu(w,c)+\mu(c,d_{c})=\mu(w,c)+n-\mu(d_{c},c)=n+1$. Hence, $w\in\operatorname{wUC}(G)$ and the only possible MS $\mathcal{X}=(\mathcal{C},\mu_{X})$ is for every candidate $c\in\mathcal{C}\setminus{\{w\}}$, $\mu(w,c)=1$, $\mu(d_{c},c)=n$ and $\mu_{X}=0$ everywhere else. $|\mathcal{X}|=(m-1)(1+n)$ and since it is the only MS, it is also an SMS. ∎
A.3 Proofs for Section 4.3
@@
Building an MS for $w\in\operatorname{MM}(G)$ is rather straightforward. We simply need to ensure that $w$ wins a minimum number $t$ of pairwise comparisons in each head-to-head with another candidate and ensure that each losing candidate loses at least ntn-t pairwise comparisons against another candidate. Suppose we do not do that and let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be an MS for $w\in\operatorname{MM}(G)$. Suppose for all $t\leq\left\lceil n/2\right\rceil$ exists $c_{0}\in\mathcal{C}\setminus{\{w\}}$ such that $\mu_{\mathcal{X}}(w,c_{0})<t$ and for all $c^{\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}}(c^{\prime},c_{0})<n-t$. Then let $\mathcal{X}^{\prime}$ be a completion of $\mathcal{X}$ such that $\mu_{\mathcal{X}^{\prime}}(w,c_{0})=\mu_{\mathcal{X}}(w,c_{0})$ and for all $c^{\prime}\in\mathcal{C}$, $\mu_{\mathcal{X}^{\prime}}(c^{\prime},c_{0})=\mu_{\mathcal{X}}(c^{\prime},c_{0})$. Then the maximin score of $w$ is $\sigma_{w}\operatorname{MM}(w,\mathcal{X}^{\prime})\geq\mu_{\mathcal{X}^{\prime}}(w,c_{0})<t$ while the maximin score of $c_{0}$ is $\sigma_{c_{0}}\operatorname{MM}(c_{0},\mathcal{X}^{\prime})\geq n-\max_{c^{\prime}\in\mathcal{C}}\mu_{\mathcal{X}}(c^{\prime},c_{0})>n-(n-t)>t$. Since $\sigma_{\operatorname{MM}}(c_{0},\mathcal{X}^{\prime})>\sigma_{\operatorname{MM}}(w,\mathcal{X}^{\prime})$, $w\not\in\text{NW}_{\operatorname{MM}}(\mathcal{X})$. Contradiction. ∎
@@
The lower bound is achieved by maximizing the values of $t$ and $k=|\{c:c\in\mathcal{C},\,\mu(w,c)\geq n-t\}|$. Typically, if for every candidate $c\in\mathcal{C}\setminus{\{w\}}$, $\mu(w,c)=n$, we have $t=\left\lfloor n/2\right\rfloor$ and k=m1k=m-1 and $|\mathcal{X}|=(n-\left\lfloor n/2\right\rfloor)(m-1)+n(m-1-(m-1)))=\left\lceil n/2\right\rceil(m-1)$. Conversely, if $k=0$, $|\mathcal{X}|=(n-t)0+n(m-1-0))=n(m-1)$. ∎
@@
Following the ideas developed in Theorem 12, we introduce Algorithm 1. The SMS returned by the algorithm maximizes the maximin score of $w$, unless its score is larger than $\left\lfloor n/2\right\rfloor$, because ensuring a larger score would break the subset minimality of the SMSs. Hence, the maxwin property is satified. ∎